### v1

# Accurate Checkerboard Corner Detection under Defoucs

Abstract
Camera calibration is a critical process in 3D vision, impacting applications in autonomous driving, robotics, architecture, and so on. This paper focuses on enhancing feature extraction for checkerboard corner detection, a key step in calibration. We analyze existing methods, highlighting their limitations and propose a novel sub-pixel refinement approach based on symmetry, which significantly improves accuracy for visible light cameras. Unlike prior symmetry based method that assume a continuous physical pattern, our approach accounts for abrupt changes in visible light camera images and defocus effects. We introduce a simplified objective function that reduces computation time and mitigates overfitting risks. Furthermore, we derive an explicit expression for the pixel value of a blurred edge, providing insights into the relationship between pixel value and center intensity. Our method demonstrates superior performance, achieving substantial accuracy improvements over existing techniques, particularly in the context of visible light camera calibration. Our code is available from https://github.com/spdfghi/Accurate-Checkerboard-Corner-Detection-under-Defoucs.git.
1 Introduction
Camera calibration is fundamental to 3D vision, with broad applications in autonomous driving, robotics, architecture, medicine, etc. Camera calibration primarily serves two purposes: first, to establish a mapping between three-dimensional homogeneous coordinates (rays originating from the camera) and pixel points, known as intrinsic calibration; and second, to calibrate the positional relationship between the camera and other cameras [16, 34] or types of sensors, such as LiDAR [24], known as extrinsic calibration. Camera calibration methods are mainly divided into those based on target based calibration and online self-calibration methods [19], depending on the input. In certain scenarios, such as aerial robots, extrinsics can undergo significant changes even with moderate shocks or during extended operations in environments with vibrations, making self-calibration important. However, since intrinsic parameters remain relatively stable even in challenging conditions [34] and target based calibration accuracy advantage over self-calibration, they dominate in most scenarios and are a necessary preliminary step for multi-sensor extrinsic calibration.
In intrinsic calibration, the projection camera model with non-linear distortion terms commonly used today was proposed by Brown [6, 8], while Zhang established an effective method for calibration based on 2D plane point mapping [33], which is widely applied. Additionally, there are calibration methods based on line distortion correction [5] and general calibration methods [18, 25] that fully consider point and line mapping. Feature extraction is the first step of calibration, which ultimately determines the accuracy and is the focus of this paper. The two main types of calibration targets are checkerboards and circles; checkerboards have the advantage of higher robustness to distortion, while circles offer higher accuracy [15, 14, 4]. The main issue with the circles pattern is that the shape of the circle under distortion can hardly be calculated. Recently, Song et al. [28] proposed an impressive analytical method to compute the centroid of a distorted circle and outperforms checkerboard-based calibration. But this method still has limitations, as extracting centroid in the image requires a larger pattern area than traditional circle fitting method, thus restricting the position of the circle’s center from the image edge, where distortion is most significant.
To improve the detection accuracy of checkerboard corner points, a series of methods have been studied. Currently, two mainstream approaches are notable. The first is assuming vector from a corner to its adjacent area pixel is orthogonal to pixel’s gradient direction, such as implemented in OpenCV [10], which achieve good sub-pixel accuracy; however, its performance is subpar under some conditions, as tested by Schöps et al. [27] and Song et al. [28] with infrared cameras. Datta et al. discovered that the assumption of OpenCV does not hold in non fronto-parallel input images [4]. Therefore they iteratively estimate the homography matrix, and then extract point after a transform to canonical fronto-parallel image, which result in a significantly improvement. The second approach involves quadratic surface fitting [20], commonly used in MATLAB. Fitting methods not only apply to regular checkerboards but have also been extended to curved surface[31]. Ha et al. [12] expanded the fitting method to patterns with three intersecting lines, and got better reprojection error than checkerboard which has two intersection lines. Schöps et al. [27] further developed Ha’s pattern by using a star-shaped calibration board with eight intersecting lines, they proposed a novel feature extraction algorithm for the pattern [26, 27], this method takes cost by pixel differences in projections of two world points that are symmetric about pattern center. With more gradient information, they got significant improvement in accuracy. They tested their method on the expanded checkerboard pattern using both RGB and infrared cameras, achieving reprojection errors that was reduced to and , respectively, compared to the OpenCV method, which is impressive.
However, we are curious why their method shows such a significant difference between the RGB and infrared cameras, and is it possible that their method is not the most effective for the RGB cameras. We carefully considered their assumptions. First, their optimization objective function implicitly assumes that the image is captured by a pinhole camera of a continuous physical pattern, and that there should not be significant distortion. Then, they assume that the image vary smoothly among neighbor pixels. Regarding the former, we believe that the assumption of a continuous physical pattern does not always hold. For the infrared cameras, since infrared radiation is based on the spontaneous emission from the thermal motion of molecules, and it is almost impossible for the surface temperature of an object to change abruptly, the assumption of a continuous spatial pattern is reasonable. However, for the RGB cameras, which captures light reflected by objects and is related to the object’s color, abrupt changes can occur within the range of a single pixel[29]. For blurry images, defocus becomes the dominant factor. Thus, we proposed a novel objection function by consider this kind of blur, experiment of visible light camera shows agree with our assumption. Moreover, our objective function is simpler with fewer parameters, which means reduced computation time and a lower risk of overfitting. We also do a math on their smooth assumption by study a blur line model, which is the convolution of 2D step function with a Gaussian kernel. The closest model to the one we used is Ref. [1], where they addressed constructing feature detector. We provided explicit expression for pixel value the first time and used it to study the variation of differences against intensity of pixel center.
As a summary, there are two main contributions of this paper:
We propose a novel sub-pixel refinement method for checkerboard corner detection based on symmetry and provide a detailed analysis of its physical assumptions. This method shows significant accuracy improvements in visible light camera compared with existing symmetry based method and mitigates overfitting risks.
For the first time, we derive an explicit expression for the pixel value (as a region integral) of an blur edge based on the convolution model of a 2D step function with a Gaussian kernel. Using this analytical result, we analyze the difference between the pixel value and the intensity at the pixel center.
2 Related Works
Camera Calibration under defocus Duda et al. [7] proposed a method based on a localized Radon transform approximated by box filters, compared to the OpenCV method, it is more robust to noise and blur, achieving significant accuracy improvements in outdoor environments. Ha et al. [11] introduced an active targets approach for calibration under defocus, they separated the checkerboard into overlapping horizontal and vertical stripes, and then applied a novel deconvolution method to these stripes. Further, they do a refraction correction of the glass screen. This method was evaluated on real data, and shows accurate camera calibration result under severe defocus.
Active target has also been applied in varies patterns, some researchers used phase information which they considered more robust in defocusing scenes than checkerboard [17, 21], and phase-shift method for camera calibration has been developed in recent years. These methods mainly contains sets up sinusoidal stripes [32] or circular patterns [3] sequentially displayed on an LCD monitor and detect sub-pixel feature points from phase map, which is calculated by phase-shifting algorithm. And with neural network for feature extraction, progress has been made in difficult tasks such as large field-of-view stereo calibration [21].
Line Detection under Defocus. The classic Canny algorithm obtains single-pixel edges through non-maximum suppression, and Devernay et al. further refines edges to sub-pixel accuracy using quadratic fitting near maxima[30]. While these well-known methods assume edge is maxima of smooth function, Ref. [29] first considers the physical scenario of a straight edge partially falling within a pixel. By analytically integrating the area segmented by the line within one pixel, it shows the discontinuity of discrete pixels will generate bias with traditional gradient computation. But they didn’t provide an analytical result considering defoucs. The calibration pattern used by Ha et al. [11] can actually be considered as lines under defocus, where they proposed an deconvolution method but also without using analytical formula. The analytical expression for line defocus was first introduced in Ref. [22] and later used in Ref. [35] for single-image defocus estimation. Bakery et al. [1] considered the pixel value rather than point value expression, which was not addressed in Ref. [22]. The line defocus model used in this paper is similar to Bakery et al., the key difference is that their paper focused on using numerical methods to construct a feature detector. They only provided an integral form while we provide explicit results.
3 Method
3.1 Camera model
We briefly review the pinhole camera and distortion model, which will be used in our experiment part. For a point in homogeneous representation of camera coordinate, the distorted point is represented as [8]
and
where
are radial distortion coefficients and are tangential distortion coefficients.
For modern, well-manufactured cameras, the shear deformation component can be ignored, and the camera model takes the following form [13]
the image coordinates can be obtained as
3.2 Line defocus model
Before introducing our algorithm, we would like to discuss pixel continuity and the effect of image blur on interpolation. In the absence of blur, traditional subpixel techniques tend to produce significant deviations near image boundaries, which is the so called partial area effect[29]. A simple example involves two horizontally adjacent pixels, and , with a vertical boundary dividing pixel into two equal area, where the intensities of the two parts are and , respectively, while the intensity of all positions in pixel is . Using linear interpolation yields an intensity of at the location between the two pixels, which differs significantly from the true intensity . Intuitively, Gaussian blur can reduce this discrepancy, and we will provide a quantitative discussion of this effect in this section.
Checkerboard feature points can be viewed as corner points where two lines intersect, we model the original image of line that captured by a noise-free pinhole camera as step function [1]
where W and B denote constant intensity of white and black region, and d represents the distance from point to the line passing through the feature center . Let denotes the angle from image x-axis to the line, one can compute through a simple cross product
Note that the two half-planes divided by the line are not identical, so is a signed quantity.
For optical camera, we need to consider the defocus factor and optical transfer function. Bakery et al. [1] modeled these effects as a Gaussian kernel function, which is a widely used [22, 35, 11] approach. 2-D Gaussian has the form
The real blur image can be calculated as a convolution of noise-free pinhole image with the kernel. Due to the symmetry of Gaussian function, we align the y-axis with direction of the line (Fig. 2(b)-(d)), which simplifies the convolution to a straightforward integration
where is the error function
implemented by most numerical program language, like C and Matlab. Using integration by parts, we can obtain the integral
and integrate again we get
From above two equations, we then can obtain the -th pixel value by performing integral over a rectangular region centered at . For the sake of brevity, we have omitted parameters from main expression and leaved the dependence to its variable and
where
and is the distance function Eq. 7 of four corner of pixel to the line, with a scaling factor
The coefficient , , and is given in Tab. 1. Similar integral form in Eq. (13) was first introduced in [1] for the purpose of constructing feature detector; however, to best of our knowledge, explicit result of this integral has never appeared in the literature aside from this paper. Compare with Eq. (9) we can find that both of them has two part, one is the average intensity, another equals intensity difference times some quantity about the distance of point to line. For pixel value, it also contains a factor with respect to and .
Fig. 3 shows the difference between Eq. (13) and Eq. (9), where letter use the center position of pixel. We randomly select the tilt angle of the line and randomly vary the position of within the range of one pixel. It can be observed that, under typical image parameters, the maximum difference between the pixel center intensity and the pixel value becomes negligible when approaches one pixel. This condition is easily satisfied during calibration.
3.3 Refinement via symmetry
For the formulas provided in the previous section, a direct approach is to perform nonlinear optimization on the pixel values. Considering that the checkerboard corners are the intersection points of two lines, for the effects of aliasing in checkerboard corner’s center, though the pattern using a display device proposed by Ha et al. [11] can be used to mitigate it. However, using a display is not the optimal choice in all calibration scenarios, and the refraction correction of display screen may introduce extra errors.
Schöps et al. developed a feature refinement method that can be applied to a multi-intersecting line pattern [27], which is a generation of checkerboard. Their method was to minimize a symmetrical function
where is the position in world coordinate centered at every pattern center, and is the homography that projects the world coordinate into image pixels. By using the Levenberg-Marquardt(LM) algorithm, they estimates the homography matrix for every feature, the result points was derived from the homography matrix. This method makes amazing result. For eight intersecting line pattern, they reduce the medium reprojection errors to less of OpenCV method with infrared camera.
However, we believe this method has two limitations: Firstly, Eq. (16) implicitly assumes that the image is captured by a pinhole camera of a continuous physical pattern, this holds for infrared camera since the surface temperature of an object can hardly change abruptly. But for visible light cameras, which captures light reflected by objects, abrupt changes can occur within the range of a single pixel[29]. In calibration, the blur in image is mainly caused by defoucs, see Fig. 4. Secondly, distortion may also leads to a poor initialization of LM algorithm. Moreover, although straight lines can still be considered straight in a small neighborhood despite distortion, which is why the chessboard method is effective. But when number of intersection lines large than two, the mapping of these lines may not be described by a homography under distortion.
Now let’s do a more physical modification of Eq. (16) for visible light camera. Within the pixel ranges used for calibration, the Gaussian of kernel function can be considered a constant. This assumption works well even with large defoucs up to several tens of pixels [12], though in that case we should consider the environment light change. We cannot use a single homography matrix to describe this mathematical relationship. A homography matrix has eight degrees of freedom and can be determined by the correspondence of four pairs of points. Considering two intersecting lines in a checkerboard pattern, a homography matrix can be computed by selecting a bounding box around any one of the blur lines. However, the pixels near the other blur line are not necessarily determined by this matrix. In fact, the other line can be at any arbitrary angle. But from Eq. (9), we can see that as long as two points are on the same side of the half-plane and equidistant from the boundary line, they have the same value. Furthermore, when another line exists, i.e., in the standard checkerboard pattern, if two points are symmetric with respect to the feature point , then they have the same value. Therefore, we can formulate the following optimization objective
Similar to [27], we randomly sample sub-pixel points and then optimized it by LM algorithm. The Jacobian is calculated by bilinear interpolation of the discrete gradient as well. Notice that Eq. (17) only requires the symmetry of kernel function, we actually do not need to know the analytical expression of .
4 Experiment
4.1 Setup
Many previous studies have achieved excellent results on synthetic data but performed a little bit poorly on real data; therefore, we directly use real data for evaluation. To date, there is no widely recognized calibration test dataset that includes the ground truth of camera parameters, so we make our own like other researchers. We primarily evaluate our results on two datasets: one captured by the infrared camera of an Intel RealSense D435 and visible light camera HIKVISION MV-SC2005AM. The latter has more noise and distortion compared to the former. The reason we chose latter is that RGB cameras may exhibit purple fringing, which could affect our assumptions. In this paper, we focus more on the impact of defocus on calibration, so we selected a camera that captures grayscale images. Besides, we want to know how distortion will affects the algorithms we compare. We evaluate accuracy using the widely used reprojection error, which projects world points into image coordinates using estimated camera parameters and then calculate the difference of extracted feature coordinates and projected positions.
We compared the performance of three methods: OpenCV, Ref. [27] and ours. Here, we are not concerned with the success rate of detection, although that is another topic worth discussing [23, 9]. We focus solely on the accuracy of sub-pixel refinement, thus using initial feature values obtained from OpenCV. Due to variations among different calibration optimization algorithms, we consistently use OpenCV’s implementation for all feature refinement methods, which initializing with Zhang’s method [33] and employing a third-order radial and second-order tangential distortion model[8] as Eqs. 1 and 2, followed by LM method up to 30 iterations.
4.2 Results
We have noticed that few studies discuss the impact of the optimization window on accuracy in refinement algorithms. Some, like Refs. [7] and [27], mention using a fixed window for their experimental data. However, we observe that, especially in cases with large distortion and perspective effects, the impact of the optimization window size on accuracy can exceed the differences between algorithms. In Fig. 5, we show the variation of reprojection errors with optimization window size for different algorithms. Fig. 5(a) presents the results for an infrared camera, as shown in Tab. 2, the distortion of this camera is relatively small. Therefore, all algorithms achieve the best reprojection error over a large window size range. In the range tested in our experiments, methods of this paper and Ref. [27] significantly outperform OpenCV’s method. The reprojection error for OpenCV stabilizes as the window size increases, while both methods of this paper and Ref. [27] show a slight decrease from to , followed by a slow rise, with Ref. [27]’s method rising more slowly. As seen in Tab. 2, for the optimal reprojection error, our method is similar to Ref. [27]. Compared to OpenCV, our method reduces the mean and median reprojection errors by and , respectively.
Fig. 5(b) shows the results for visible light camera. Compared to the infrared camera we used, it has a lower resolution and a wider field of view. As seen in Tab. 2, it has larger camera distortion. In this case, the accuracy of both the OpenCV and Ref. [27]’s methods fluctuates significantly with the change in window size. In particular, Ref. [27]’s method shows an overall decrease, but with a fluctuation around size . We believe this is mainly due to the inappropriate physical assumptions of this method in this situation, as discussed earlier. Additionally, we observed some sample points fell outside the image boundaries during the algorithm’s execution, which is caused by difficulties in initializing the homography matrix due to distortion and noise, while the LM algorithm is sensitive to initial values. We discarded these points. However, the same algorithm did not encounter this issue in any of the windows for the infrared camera. For OpenCV, its overall error decreases first and then increases, achieving good reprojection error only in the narrow window size range of . When the window size increases to , Ref. [27]’s method reduces the error by almost compared to OpenCV. For the optimal error, we significantly outperform both methods, reducing the mean and median reprojection errors by and compared to Ref, and by and compared to OpenCV.
| Reprojection error | Camera matrix | Distortion() | |||||||||
| Methods | Average | Medium | |||||||||
| Infrared camera | |||||||||||
| OpenCV | 0.0842 | 0.0635 | 640.55 | 642.15 | 643.42 | 400.23 | 4.82 | -10.00 | 5.85 | 0.07 | 0.35 |
| Proposed | 0.0576 | 0.0430 | 640.75 | 642.34 | 643.54 | 400.31 | 4.30 | -7.55 | 4.15 | 0.14 | 0.42 |
| Ref. [27] | 0.0561 | 0.0420 | 640.80 | 642.41 | 643.52 | 400.33 | 4.34 | -7.60 | 4.19 | 0.16 | 0.40 |
| Visible light camera | |||||||||||
| OpenCV | 0.0848 | 0.0627 | 471.07 | 470.93 | 386.14 | 296.41 | 27.53 | -74.05 | 16.00 | -0.19 | -0.12 |
| Proposed | 0.0636 | 0.0470 | 471.30 | 471.17 | 386.24 | 296.43 | 24.88 | -68.17 | 11.76 | -0.19 | -0.09 |
| Ref. [27] | 0.0870 | 0.0577 | 471.29 | 471.16 | 386.04 | 296.47 | 27.71 | -76.89 | 16.34 | -0.17 | -0.29 |
Tab. 2 presents the estimated parameters, distortion coefficients are from Eqs. 1 and 2. For the infrared camera, our results are similar to those of the Ref. [27]. For the visible light camera, the camera matrix parameters estimated by the Ref. [27] are similar to ours, while the distortion parameters align more closely with those of OpenCV. Fig. 6 shows the distribution of the reprojection error. In the visible light camera case, the reference method produces large errors in several views. In the blue-colored view, outliers are observed in our method, which corresponds to the viewpoint with the highest reprojection error using the OpenCV method. The primary cause is that, in this view, some feature points are significantly farther from the camera’s z-axis, leading to higher noise levels. While this issue can be solved through such as manual correction by MATLAB toolbox, we decided to keep the original outcomes of algorithm.
5 Discussions
Conclusion In this paper, we proposed a novel feature extraction method for camera calibration, addressing the limitations of existing approaches, particularly for visible light camera cameras under defocus and distortion conditions. Our objective function, which accounts for blur nature of optical defocus, simplifies computation and reduces the risk of overfitting for symmetry based method proposed by Schöps et al.. Through experiments on real-world data from both infrared and visible light camera cameras, we demonstrated that our method significantly improves reprojection error compared to OpenCV and approach by Schöps et al.. Specifically, we achieved a mean and medium reprojection error reduction of and than Schöps et al. for visible light camera cameras with large distortion. For infrared case, we get a similar result with optimal refinement window size. In both types of cameras, we achieved significantly better results compared to OpenCV, reduce reprojection error to . We also discussed the impact of optimization window size on the algorithms and found that it is important for cameras with large distortion. In such cases, OpenCV is highly sensitive to the optimization window size, while our method is more robust.
We discussed two different causes of blurred images and explained why the method by Schöps et al. is not suitable for visible light cameras. For visible light cameras, we considered the Gaussian point spread function blur model for straight edges of checkerboard patterns and provided an analytical solution for discrete pixel values. This is significant for further calibration methods based on deconvolution, as well as precise localization in applications like QR code detection. Our calculations show that when defocus is small, there is a significant difference between point intensity and pixel intensity, which may lead to more accurate results in current single-image depth from defocus methods based on point intensity.
In conclusion, our method offers a significant improvement in feature extraction accuracy for camera calibration, especially in real-world scenarios where distortion, defocus, and challenging poses are present. These findings suggest that our approach is well-suited for practical applications in 3D vision, autonomous systems, and multi-sensor environments, where reliable and accurate calibration is critical.
Limitations and feature works For the visible light camera, we calibrated not only industrial cameras but also smartphone cameras. Although our method still achieved better results than OpenCV, the difference was not as significant as we expected. We observed noticeable purple fringing in the smartphone camera when capturing the checkerboard pattern, which may be caused by chromatic aberration. Therefore, there is room for algorithmic improvement for cheap digital cameras. Another issue is, although our experiments show that the method by Schöps et al. is not optimal for visible light cameras, it still achieved significantly better results than OpenCV with the star pattern they proposed. This is because, as explained by Datta et al., the assumptions of the OpenCV method only hold for non fronto-parallel input images, meaning at most two intersecting lines meet the assumptions, which limits its ability to effectively utilize the information from all lines. However, as reported by Schöps et al. increasing the number of intersection lines on the star pattern beyond does not further enhance accuracy. How to resolve aliasing at the center of star based pattern to achieve higher precision remains to be studied further. This will be a direction for our future work.